OpenStudy (anonymous):

0.01 = A. 1–1 B.1–2 C.10–1 D.10–2

OpenStudy (anonymous):

help

OpenStudy (anonymous):

i guess .01 =1/100 now 100 =10^2 so 1/100 = 10^(-2) so D.10–2

OpenStudy (anonymous):

0.01 = 1^–1 1^–2 10^–1 10^–2

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Which is NOT a power of 10? 10^-6 0 1 100

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

0

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

R U SURE

OpenStudy (anonymous):

HELLO

OpenStudy (anonymous):

if 10 is raised to some power ,then does the result is 0

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

welcome bye